# MolMix: A Simple Yet Effective Baseline for Multimodal Molecular Representation Learning

### v1

@@ 3 Experimental Setup and Results
Table 3: Modality ablation study on the Kraken dataset (MAE $\downarrow$). We keep the downstream Transformer fixed and train using a single modality or a combination of modalities. Using all three modalities obtains the best results on three out of the four properties, with the second-best results generally being obtained by a configuration that contains 3D conformers. Notably, for the buried Sterimol L property, the best results are obtained by a 3D encoder + Transformer model, indicating that the property could mainly depend on the 3D structure.
@@
For all experiments, we report the mean and standard deviation over five different runs. We use the same hyperparameters for the modality encoders—we encode the SMILES strings using a transformer with two layers and four attention heads, the 2D graph is processed by a GINE [29] network containing 6 layers, and the 3D conformations are processed either using a SchNet [8] or a GemNet [9] model. All of the encoders have a hidden dimension of 128. The modality embeddings are then projected by a linear layer to a 512-dimensional space before we jointly processing them with a downstream Transformer network with 8 heads and 6 layers for the Lipo, ESOL, FreeSolv, BACE and Kraken, and 12 heads with 8 layers on Drugs-75K.
Table 4: Transfer learning experiment. We select the best checkpoint of a model trained to predict the electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze the model and only train the last linear readout layer on the Kraken dataset. We compare with a randomly initialized model. For all descriptors, using the pre-trained weights improve predictive performance. Note that pretraining improves both mean performance and standard deviations.
@@
To address Question 2, table 3 shows that for three of the four Kraken descriptors, training on all three modalities yields the best results, while for one property, training on the 3D modality alone performs slightly better. The results indicate that 3D tokens contribute the most to downstream performance, followed by 2D, with SMILES (1D) having the least impact. Notably, SMILES strings significantly improve performance when predicting the Sterimol L descriptor.
To answer Question 3, we select the best checkpoint from a model trained to predict electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze its weights and train only the final linear layer to predict descriptors on the Kraken dataset, comparing it to a randomly-initialized model. As shown in table 4, pre-training improves predictive performance in all cases, suggesting that with sufficient data, MolMix could serve as a foundation model for molecular tasks.

### v2

@@ 3 Experimental Setup and Results
Figure 1: Modality ablation study on the Kraken dataset (MAE $\downarrow$). We keep the downstream Transformer fixed and train using a single modality or a combination of modalities. Using all three modalities obtains the best results on three out of the four properties, with the second-best results generally being obtained by a configuration that contains 3D conformers. Notably, for the buried Sterimol L property, the best results are obtained by a 3D encoder + Transformer model, indicating that the property could mainly depend on the 3D structure.
@@
Figure 2: Transfer learning experiment. We select the best checkpoint of a model trained to predict the electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze the model and only train the last linear readout layer on the Kraken dataset. We compare with a randomly initialized model. For all descriptors, using the pre-trained weights improve predictive performance. Note that pretraining improves both mean performance and standard deviations.
@@
For all experiments, we report the mean and standard deviation over five different runs. We use the same hyperparameters for the modality encoders—we encode the SMILES strings using a transformer with two layers and four attention heads, the 2D graph is processed by a GINE [29] network containing 6 layers, and the 3D conformations are processed either using a SchNet [8] or a GemNet [9] model. All of the encoders have a hidden dimension of 128. The modality embeddings are then projected by a linear layer to a 512-dimensional space before we jointly processing them with a downstream Transformer network with 8 heads and 6 layers for the Lipo, ESOL, FreeSolv, BACE and Kraken, and 12 heads with 8 layers on Drugs-75K. We use the Schedule-Free AdamW optimizer [36, 37, 38].
@@
To address Question 2, fig. 2 shows that for three of the four Kraken descriptors, training on all three modalities yields the best results, while for one property, training on the 3D modality alone performs slightly better. The results indicate that 3D tokens contribute the most to downstream performance, followed by 2D, with SMILES (1D) having the least impact. Notably, SMILES strings significantly improve performance when predicting the Sterimol L descriptor.
To answer Question 3, we select the best checkpoint from a model trained to predict electronegativity ($\chi$) on the Drugs-75K dataset. We then freeze its weights and train only the final linear layer to predict descriptors on the Kraken dataset, comparing it to a randomly-initialized model. As shown in fig. 2, pre-training improves predictive performance in all cases, suggesting that with sufficient data, MolMix could serve as a foundation model for molecular tasks.